Chris Hanusa

DISCLAIMER

These “Answer Keys” is a presentation of what a homework assignment should look like, insofar as to answer each question asked by the homework assignment. This will not be as complete as an answer that receives the maximum number of points on an assignment; it is just here for you to get an idea of what I was looking for while grading. There will be many answers that work as well as mine.

It will be rare that I change grades on homework. But if you wish to come and talk with me about my grading or want help with the next homework assignment, I am available during my office hours, Wednesday mornings 10-11am or by appointment. E-mail me at chanusa@washington.edu.